OpenStudy (anonymous):

Can anyone explain what "COMBINING LIKE TERMS" is and what each thing that needs to be recognized is? (Tell me what the constants, like terms, and other things that need to be recognized is..) PLEASE. and explain it in distributive property!!

OpenStudy (anonymous):

9a+4a, the a's are like terms, so you add them together, thus, getting 13a unless it's something like 4a+9a=5a(9+5) then you only add the like terms that are on the same side of the equal sign.

OpenStudy (anonymous):

Basically you want to perform arithmetic on numbers that have the same letter ie. 10x+4x-3z+7z-2t 10x+4x=14x -3z+7z=4z -2t=-2t Distributive property involves parenthesis. multiply the number on the outside of the parenthesis by the 2 numbers inside or how ever many terms you have inside 2(2x-4) so you would multiply 2 by 2x and 2 by -4 2(2x-4)=4x-8 also familiarize your self with the term FOIL for future math classes First Outer Inner Last